OpenStudy (anonymous):

Please help. This is summer work for AP Calc. SIMPLIFY ** (1/x+h) - (1/x) **

OpenStudy (dumbcow):

this for finding derivative of 1/x using limits right? \[\lim_{h \rightarrow 0} \frac{\frac{1}{x+h} - \frac{1}{x}}{h}\]

OpenStudy (dumbcow):

combine fractions using common denominator "x(x+h)" \[\rightarrow \frac{x - (x+h)}{x (x+h)}\]

OpenStudy (anonymous):

Oh. Thank you!

OpenStudy (dumbcow):

no problem :)

OpenStudy (anonymous):

Wait. Would that be the simplest form? Or would any terms cancel out?

OpenStudy (dumbcow):

you tell me ... what happens when you distribute the negative and add x-(x+h) = x -x -h

OpenStudy (anonymous):

It would be (-h) / x(x+h).. Right?

OpenStudy (dumbcow):

yep

OpenStudy (anonymous):

Thank you so much